OpenStudy (anonymous):

Rewrite the equation below in standard form -6x+3y=12y

OpenStudy (anonymous):

please...

OpenStudy (anonymous):

a.3x-6y=12 b.3x-6y=-12 c.-x+2y=4 d.x-2y=-4

OpenStudy (anonymous):

those are my options

jigglypuff314 (jigglypuff314):

Hello and Welcome to Openstudy! :) could you try revising your question? there might have been a typo :) I see: -6x +3y = 12y

OpenStudy (anonymous):

-3x+6y=12 sorry:(

jigglypuff314 (jigglypuff314):

much better :) now tell me, what is the Greatest Common Factor of 3, 6, and 12?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

i think

jigglypuff314 (jigglypuff314):

yep :) so you can divide all the values by 3. -3x divided by 3 = -x 6y divided by 3 = 2y 12 divided by 3 = 4

OpenStudy (anonymous):

omg thank you so much :)

jigglypuff314 (jigglypuff314):

you're welcome :) glad I could help

OpenStudy (anonymous):

Use substitution to determine which of the following points is a solution to the standard form equation below. A. (10,5/2) B. (5,2/5) C. (5/2,-5/2) D. (5/2,-2/5)

OpenStudy (anonymous):

can you help me with one more please?

jigglypuff314 (jigglypuff314):

sure :) what's the given equation that they want you to use? the way you would do it would be to plug in the coordinate pairs as (x, y) so for point (10, 5/2) you would plug in 10 for x and 5/2 for y then simplify and if the two sides equal each other then it works :)

OpenStudy (anonymous):

2y=5 is the equation

jigglypuff314 (jigglypuff314):

could you check again? perhaps you made a typo :)

OpenStudy (anonymous):

uumm it says 2y=5 unless the paper is wrong :o

jigglypuff314 (jigglypuff314):

weird... I suppose it would still work ;) could you solve for y given 2y = 5 ? just divide both sides by 2

OpenStudy (anonymous):

ok it worked thank you so much :)

jigglypuff314 (jigglypuff314):

you're welcome :)